# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ Acknowledgement
$D_{\beta,P}^{r}=I_{P}^{-1}(A-g_{P}(W)){\bf\phi}(W)(Y-\tilde{\theta}_{P}^{r}(W)-(A-g_{P}(W))\sum_{j}\beta_{P}(j)\phi_{j}(W)),$
where $I_{P}=E_{P}g(1-g)(W){\bf\phi}{\bf\phi}^{\top}$. ∎
Appendix B: Data-Generating Processes, Missing Outcome Data Structure and Simulations
The data-generating process for the SBP synthetic study in Section 6 is:
$U_{Y}$ $\sim\mathcal{N}(0,1)$
$\text{Age}=W_{1}$ $\sim\mathcal{N}(50,12^{2})$
$\text{BMI}=W_{2}$ $\sim\mathcal{N}(28,5^{2})$
$\text{Smoking}=W_{3}$ $\sim\text{Bern}(0.2)$
@@
$A$$\sim\left\{\begin{array}[]{ll}\text{Bern}(0.5)&\text{if }S=1\\ \text{Bern}(\text{expit}(-1.5+0.05W_{1}+0.04W_{2}+0.8W_{3}+0.03W_{4}+0.02W_{5}))&\text{if }S=0\end{array}\right.$
$Y$ $=120-0.1W_{1}+0.05W_{2}^{2}-5W_{3}+0.04W_{4}+0.03W_{5}+U_{Y}.$
@@
$B$ =$=\left\{\begin{array}[]{ll}0.2+0.1W_{1}\cdot I(A=0)&\text{scenario (a)}\\ 0.5+3.1W_{1}\cdot I(A=0)+0.8W_{3}&\text{scenario (b)}\end{array}\right.$
@@
The data-generating process for scenario (e) in Section 8 is:
$U_{Y}$ $\sim\mathcal{N}(0,1)$
$W_{1}$ $\sim\text{Uniform}(0,1)$
$W_{2}$ $\sim\text{Uniform}(0,1)$
$W_{3}$ $\sim\text{Bern}(0.5)$
$A$ $\sim\text{Bern}(0.67)$
$B$ $=0.2+0.1W_{1}\cdot I(A=0)$
$Y$ $\sim\text{Bern}(\text{expit}(-3-0.9W_{1}+0.5W_{2}-0.7W_{3}+0.9A+I(S=0)\cdot B)).$
@@
$\Psi(P_{0})=\mathbb{E}_{0}[\mathbb{E}_{0}(Y\mid S=1,W,A=1,\Delta=1)-\mathbb{E}_{0}(Y\mid S=1,W,A=0,\Delta=1)].$
Now, to learn the working models for the effect of $S$ on $Y$ and the effect of $A$ on $Y$, we use the following two loss functions, respectively:
$L_{\tilde{\theta}_{P},\Pi_{P},g^{\Delta}_{P}}(\tau)$ $=\frac{\Delta}{g^{\Delta}_{P}(1\mid S,W,A)}\left(Y-\tilde{\theta}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))\tau\right)^{2},$
$L_{\tilde{\theta}^{r}_{P},g_{P},\tilde{g}^{\Delta}_{P}}(T)$ $=\frac{\Delta}{\tilde{g}^{\Delta}_{P}(1\mid W,A)}(Y-\tilde{\theta}^{r}_{P}(W)-(A-g_{P}(1\mid W))T(W))^{2},$
where $g^{\Delta}_{P}(1\mid S,W,A)\equiv P(\Delta=1\mid S,W,A)$ and $\tilde{g}^{\Delta}_{P}(1\mid W,A)\equiv P(\Delta=1\mid W,A)$. In addition, we also multiply the factor $\Delta/g^{\Delta}_{P}(1\mid S,W,A)$ and $\Delta/\tilde{g}^{\Delta}_{P}(1\mid W,A)$ to the $Y$-components of the canonical gradients of the projection parameters we showed in Lemma 5 and 8 respectively. We conducted simulations for the scenario where outcome is subject to missingness. The data-generating process is:
$U_{Y}$ $\sim\mathcal{N}(0,1)$
$W_{1}$ $\sim\mathcal{N}(0,1)$
$W_{2}$ $\sim\mathcal{N}(0,1)$
$W_{3}$ $\sim\mathcal{N}(0,1)$
$B$ =$=\left\{\begin{array}[]{ll}0.2+0.1W_{1}\cdot I(A=0)&\text{scenario (a)}\\ 0.5+3.1W_{1}\cdot I(A=0)+0.8W_{3}&\text{scenario (b)}\end{array}\right.$
$\Delta$ $=\text{Bern}(\text{expit}(2.3+0.8W_{1}))$
$Y$ $=\Delta[2.5+0.9W_{1}+1.1W_{2}+2.7W_{3}+1.5A+U_{Y}+I(S=0)\cdot B].$
See figure 5 for results.
Figure 5: Comparing the MSE, relative MSE, 95% confidence interval coverage of A-TMLE, ES-CVTMLE, and TMLE estimator for increasing sample sizes of both the RCT and external data. The reference (dashed red line) in the relative MSE plots are the MSEs of the RCT-only estimator. The plots in the right-most column show the average proportion of cross-validation folds the ES-CVTMLE estimator pooled the external data.
@@
The A-TMLE is targeting a data dependent target parameter $\Psi_{{\cal M}_{w,n}}(P_{0})$ ignoring that the parameter itself already depends on the data. In the literature for data-dependent target parameters, we have generally recommended the use of cross-validated TMLE (CV-TMLE) to minimize reliance on Donsker class conditions. Therefore, in this appendix, we describe and analyze the adaptive CV-TMLE for this particular type of data-adaptive target parameter $P_{n}\rightarrow\Psi_{{\cal M}_{w}(P_{n})}(P_{0})$.
Let $P_{n,v}$ and $P_{n,v}^{1}$ be the empirical measures of the training and validation sample, respectively, for $v=1,\ldots,V$, using $V$-fold cross-validation scheme. Let ${\cal M}_{w,n,v}={\cal M}_{w}(P_{n,v})$. This defines a data-adaptive target parameter $\Psi_{{\cal M}_{w,n,v}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$. Let $P_{n,v}^{0}$ be an initial estimator of $P_{0}$ based on $P_{n,v}$. Let $P_{n,v}^{*}$ be a TMLE of $\Psi_{{\cal M}_{w,n,v}}(P_{0})$ based on this initial estimator $P_{n,v}^{0}$, where the TMLE-update is applied to the validation sample $P_{n,v}^{1}$, so that $P_{n,v}^{1}D_{{\cal M}_{w,n,v},P_{n,v}^{*}}=o_{P}(n^{-1/2})$, for $v=1,\ldots,V$. One could also carry out a pooled-TMLE-update step by minimizing the cross-validated empirical risk $\epsilon\rightarrow 1/V\sum_{v}P_{n,v}^{1}L(P_{n,v}^{0}(\epsilon))$ using the same least favorable path $\{P_{n,v}^{0}(\epsilon):\epsilon\}$. In both cases, we end up with $P_{n,v}^{*}$, $v=1,\ldots,V$, that solves $1/V\sum_{v=1}^{V}P_{n,v}^{1}D_{{\cal M}_{n,w,v},P_{n,v}^{*}}=o_{P}(n^{-1/2})$. We note that if ${\cal M}_{w,n,v}$ is a parametric model, and a sample size of $n/V$ is large enough to train the parameters of this working model, then one could estimate $P_{0}$ with an MLE $P_{n,v}^{*}=\arg\min_{P\in{\cal M}_{w,n,v}}P_{n,v}^{1}L(P)$, thus not using an initial estimator. Using a least favorable path with many more extra parameters in the pooled-TMLE allows one to use a simple initial estimator (less data-adaptive), the pooled-TMLE starts resembling a parametric MLE over the cross-validated empirical risk.
@@
$\psi_{n}^{*}\equiv\frac{1}{V}\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n,v}}(P_{n,v}^{*}).$
@@
$\psi_{n}^{*}-\frac{1}{V}\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n,v}}(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{*}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{*},P_{0})+o_{P}(n^{-1/2}).$
We assume the analogue of $\Psi_{{\cal M}_{w,n}}(P_{0})-\Psi(P_{0})=o_{P}(n^{-1/2})$ as presented in in Section 7, which is now given by
@@
$\psi_{n}^{*}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{*}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{*},P_{0})+o_{P}(n^{-1/2}).$
The analogue rate of convergence condition is given by $1/V\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{*},P_{0})=o_{P}(n^{-1/2})$. If conditional on the training sample $P_{n,v}$, $D_{{\cal M}_{w,n,v},P_{n,v}^{*}}$ falls in a $P_{0}$-Donsker class (note that only the targeting step depends on $P_{n,v}^{1}$, making this a very weak condition), then it follows
$\psi_{n}^{*}-\Psi(P_{0})=O_{P}(n^{-1/2}).$
Suppose $\parallel D_{{\cal M}_{w,n,v},P_{n,v}^{*}}-D_{{\cal M}_{w,n,v},P_{n,v,0}^{*}}\parallel_{P_{0}}=o_{P}(1)$, where $P_{n,v,0}^{*}$ is the TMLE-update of $P_{n,v}^{0}$ under $P_{0}$ (i.e., maximize $P_{0}L(P_{n,v,\epsilon}^{0})$ instead of $P_{n,v}^{1}L(P_{n,v,\epsilon}^{0})$ for a universal least favorable path through $P_{n,v}^{0}$). Then, we obtain
$\psi_{n}^{*}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v,0}^{*}}+o_{P}(n^{-1/2}).$
For each $v$, we have that $(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v,0}^{*}}$ is a sum of mean zero independent random variables so that this term will be asymptotically normal if the variance converges to a fixed $\sigma^{2}_{0}$. Let’s assume that $\parallel D_{{\cal M}_{w,n,v},P_{n,v,0}^{*}}-D_{{\cal M}_{w,n,v},P_{0}}\parallel_{P_{0}}=o_{P}(1)$ so that we obtain
$\psi_{n}^{*}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{0}}+o_{P}(n^{-1/2}).$
@@
$\sigma_{w,n}^{-1}(\psi_{n}^{*}-\psi_{0})=\sigma_{w,n}^{-1}\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{0}}/\sigma_{w,n,v}+o_{P}(n^{-1/2}).$
@@
Assume $1/V\sum_{v}P_{n,v}^{1}D_{{\cal M}_{w,n,v},P_{n,v}^{*}}=o_{P}(n^{-1/2})$. We have
$\psi_{n}^{*}-\frac{1}{V}\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n,v}}(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{*}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{*},P_{0})+o_{P}(n^{-1/2}).$
@@
$\psi_{n}^{*}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{*}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{*},P_{0})+o_{P}(n^{-1/2}).$
Rate of convergence condition: Assume $1/V\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{*},P_{0})=o_{P}(n^{-1/2})$.
Weak Donsker class condition: Assume, conditional on the training sample $P_{n,v}$, $D_{{\cal M}_{w,n,v},P_{n,v}^{*}}$ falls in a $P_{0}$-Donsker class (note that only the targeting step depends on $P_{n,v}^{1}$, making this a very weak condition). Then,
$\psi_{n}^{*}-\Psi(P_{0})=O_{P}(n^{-1/2}).$
Consistency of TMLE $P_{n,v}^{*}$ to $P_{0}$: Assume that $\parallel D_{{\cal M}_{w,n,v},P_{n,v,0}^{*}}-D_{{\cal M}_{w,n,v},P_{0}}\parallel_{P_{0}}=o_{P}(1)$. Let $\sigma_{w,n,v}^{2}=P_{0}D_{{\cal M}_{w,n,v},P_{0}}^{2}$ and $\sigma_{w,n}=1/V\sum_{v=1}^{V}\sigma_{w,n,v}$. Then,
$\sigma_{w,n}^{-1}(\psi_{n}^{*}-\psi_{0})=\sigma_{w,n}^{-1}\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{0}}/\sigma_{w,n,v}+o_{P}(n^{-1/2}).$
@@
Meta-HAL-MLE to generate working model for $\tau_{0}$: To obtain extra signal in the data for the second regression targeting $\tau_{0}$ we might apply cross-fitting again, analogue to above for the parametric regression working model. Let $P_{n,v}$ and $P_{n,v}^{1}$ be the training and validation sample for the $v$-th sample split, respectively, and $\theta_{n,v}=\hat{\Theta}(P_{n,v})$ be the resulting training sample fit of $\theta_{0}$, $v=1,\ldots,V$. We can define $\beta_{n}^{C}=\arg\min_{\beta,\parallel\beta\parallel\leq C}1/V\sum_{v=1}^{V}P_{n,v}^{1}L(\theta_{n,v}+\sum_{j}\beta(j)S\phi_{j})$ as the minimizer of the cross-fitted empirical risk. We can select $C$ with cross-validation resulting in a $\beta_{n}^{cf}$ which then generates the working model $\{\sum_{j,\beta_{n}^{cf}(j)\not=0}S\beta(j)\phi_{j}:\beta\}$ for $\tau_{0}$. One can think of $\theta_{n}+\sum_{j}\beta_{n}^{cf}(j)S\phi_{j}$ as a meta-HAL-MLE, where cross-validation selects the best $\beta$-specific estimator $\hat{\theta}(P_{n})+\sum_{j}S\beta(j)\phi_{j}$, under an $L_{1}$-norm constraint on $\beta$. Carrying out this meta-HAL-MLE (using cross-validation to select $C$) implies the semiparametric working model ${\cal Q}_{w,sp,n}=\{\theta+\sum_{j,\beta_{n}^{cf}(j)\not=0}\beta(j)S\phi_{j}:\beta,\theta\}$.
@@
Imagine that one has access to a previous related study that collected the same covariates and outcome with possibly a different treatment. One could use this previous study to learn a working model for the outcome regression $\bar{Q}_{0}$, possibly apply some undersmoothing to make it not too adaptive towards the true regression function in that previous study. One wants to make sure that that study has a sample size larger or equal than the one in the current study so that the resulting working model is flexible enough for the sample size of the current study. One could now use this ${\cal M}_{w,n}$ as working model in the A-TMLE. One can now either use a TMLE or CV-TMLE of the target parameter $\Psi_{{\cal M}_{w,n}}(P_{0})$. Below we present the theorem for the resulting adaptive CV-TMLE. The advantage of learning the working model on an external data set is that it allows us to establish asymptotic normality without requiring that $D_{{\cal M}_{w,n},P_{0}}$ converges to a fixed $D_{{\cal M}_{0},P_{0}}$. One still needs that ${\cal M}_{w,n}$ approximates $P_{0}$ in the sense that $\Psi_{{\cal M}_{w,n}}(P_{0})-\Psi(P_{0})=o_{P}(n^{-1/2})$, but as we argued in [adml] that condition can be achieved without relying on ${\cal M}_{w,n}$ to converge to a fixed oracle model ${\cal M}_{0}$ (in essence only relying on ${\cal M}_{w,n}$ to approximately capture $P_{0}$).

### v2

@@ Acknowledgement
$D_{\beta,P}^{r}=I_{P}^{-1}(A-g_{P}(1\mid W)){\bf\phi}(W)(Y-\theta_{P}(W)-(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)),$
where $I_{P}=E_{P}g(1-g)(1\mid W){\bf\phi}{\bf\phi}^{\top}$. ∎
Proof of Lemma 6.
The proof is analogous to the proof of Lemma 3 by replacing $A$ with $S$. Specifically, in Lemma 3, we showed the loss function for learning the working model for the conditional effect of $A$ on $Y$. Here, in Lemma 6, we have a similar loss function for learning the conditional effect of $S$ on $Y$. ∎
Proof of lemma 7.
Proof of lemma 8.
The $p_{Y\mid S,W,A}$-score component of $D_{{\cal M}_{S,w},P}^{\#}$ is given by
$D_{{\cal M}_{S,w},\beta,P}^{\#}$$\equiv\left\{E_{P}\Pi_{P}(0\mid W,0)\frac{d}{d\beta_{P}}\tau_{S,\beta_{P}}(W,0)\right\}^{\top}D_{\beta,P}$
$-\left\{E_{P}\Pi_{P}(0\mid W,1)\frac{d}{d\beta_{P}}\tau_{S,\beta_{P}}(W,1)\right\}^{\top}D_{\beta,P}$
The $p_{W}$-score component of $D_{{\cal M}_{S,w},P}^{\#}$ is given by
$D^{\#}_{{\cal M}_{S,w},P_{W},P}\equiv\Pi_{P}(0\mid W,0)\tau_{S,\beta_{P}}(W,0)-\Pi_{P}(0\mid W,1)\tau_{S,\beta_{P}}(W,1)-\Psi_{{\cal M}_{S,w}}^{\#}(P).$
Finally, we need the contribution from the dependence of $\Psi^{\#}_{{\cal M}_{S,w}}(P)$ on the conditional distribution $\Pi_{P}$. The influence curve of $\Pi_{P}(0\mid w,a)$ is given by $D_{{\cal M}_{S,w},\Pi,(w,a),P}=I(W=w,A=a)/P(w,a)(I(S=0)-\Pi_{P}(0\mid w,a))$. So, $\Pi_{P}$-score component of $D_{{\cal M}_{S,w},P}^{\#}$ is given by
@@
=$=\left\{\frac{I(A=0)}{g_{P}(0\mid W)}\tau_{S,\beta_{P}}(W,0)-\frac{I(A=1)}{g_{P}(1\mid W)}\tau_{S,\beta_{P}}(W,1)\right\}(I(S=0)-\Pi_{P}(0\mid W,A)).$
Now, note that $I(S=0)-\Pi(0\mid W,A)=-(I(S=1)-\Pi(1\mid W,A))$. Therefore, we have
$D^{\#}_{{\cal M}_{S,w},\Pi,P}=\left\{\frac{A}{g_{P}(1\mid W)}\tau_{S,\beta_{P}}(W,0)-\frac{1-A}{g_{P}(0\mid W)}\tau_{S,\beta_{P}}(W,1)\right\}(S-\Pi_{P}(1\mid W,A)).$
So, we have every component in
$D^{\#}_{{\cal M}_{S,w},P}=D^{\#}_{{\cal M}_{S,w},P_{W},P}+D^{\#}_{{\cal M}_{S,w},\Pi,P}+D^{\#}_{{\cal M}_{S,w},\beta,P}.$
Appendix B: Data-Generating Processes, Missing Outcome Data Structure
@@
$B$ =$=\left\{\begin{array}[]{ll}0.2+1.1W_{1}\cdot I(A=0)&\text{scenario (a)}\\ 0.5+3.1W_{1}\cdot I(A=0)+0.8W_{3}&\text{scenario (b)}\end{array}\right.$
@@
The data-generating process for scenarios with practical violations of trial enrollment positivity in Section 8 is:
$U_{Y}$ $\sim\mathcal{N}(0,0.2^{2})$
$W_{1}$ $\sim\text{Uniform}(-1,1)$
$W_{2}$ $\sim\text{Uniform}(-1,1)$
$W_{3}$ $\sim\text{Uniform}(-1,1)$
$S$ $\sim\text{Bern}(\text{expit}(\alpha(-2+W_{1}+W_{2}+\sin(2W_{1})+\sin(2W_{2})))),\ \alpha\in\{0.5,1\}$
$A$$\sim\left\{\begin{array}[]{ll}\text{Bern}(0.67)&\text{if }S=1\\ \text{Bern}(\text{expit}(-0.5W_{1}))&\text{if }S=0\end{array}\right.$
$B$ $=0.2+2.1W_{1}A$
$Y$ $=1.9+1.5A+0.9W_{1}+1.4W_{2}+2.1W_{3}+U_{Y}+I(S=0)\cdot B.$
@@
$\Psi(P_{0})=E_{0}[E_{0}(Y\mid S=1,W,A=1,\Delta=1)-E_{0}(Y\mid S=1,W,A=0,\Delta=1)].$
Now, to learn the working models for the effect of $S$ on $Y$ and the effect of $A$ on $Y$, we use the following loss functions, respectively:
$L_{\bar{Q}_{P},\Pi_{P},g^{\Delta}_{P}}(\tau)$ $=\frac{\Delta}{g^{\Delta}_{P}(1\mid S,W,A)}\left(Y-\bar{Q}_{P}(W,A)-(S-\Pi_{P}(1\mid W,A))\tau\right)^{2},$
$L_{\theta_{P},g_{P},\tilde{g}^{\Delta}_{P}}(\tau)$ $=\frac{\Delta}{\tilde{g}^{\Delta}_{P}(1\mid W,A)}\left(Y-\theta_{P}(W)-(A-g_{P}(1\mid W))\tau\right)^{2},$
where $g^{\Delta}_{P}(1\mid S,W,A)\equiv P(\Delta=1\mid S,W,A)$ and $\tilde{g}^{\Delta}_{P}(1\mid W,A)\equiv P(\Delta=1\mid W,A)$. In addition, we also multiply the factor $\Delta/g^{\Delta}_{P}(1\mid S,W,A)$ and $\Delta/\tilde{g}^{\Delta}_{P}(1\mid W,A)$ to the $Y$-components of the canonical gradients of the projection parameters we showed in Lemma 5 and 8 respectively.
@@
The A-TMLE is targeting a data dependent target parameter $\Psi_{{\cal M}_{w}}(P_{0})$ ignoring that the parameter itself already depends on the data. In the literature for data-dependent target parameters, we have generally recommended the use of cross-validated TMLE (CV-TMLE) to minimize reliance on Donsker class conditions. Therefore, in this appendix, we describe and analyze the adaptive CV-TMLE for this particular type of data-adaptive target parameter $P_{n}\rightarrow\Psi_{{\cal M}_{w}(P_{n})}(P_{0})$.
Let $P_{n,v}$ and $P_{n,v}^{1}$ be the empirical measures of the training and validation sample, respectively, for $v=1,\ldots,V$, using $V$-fold cross-validation scheme. Let ${\cal M}_{w,n,v}={\cal M}_{w}(P_{n,v})$. This defines a data-adaptive target parameter $\Psi_{{\cal M}_{w,n,v}}:{\cal M}\rightarrow\hbox{${\rm I\kern-1.99997ptR}$}$. Let $P_{n,v}^{0}$ be an initial estimator of $P_{0}$ based on $P_{n,v}$. Let $P_{n,v}^{\star}$ be a TMLE of $\Psi_{{\cal M}_{w,n,v}}(P_{0})$ based on this initial estimator $P_{n,v}^{0}$, where the TMLE-update is applied to the validation sample $P_{n,v}^{1}$, so that $P_{n,v}^{1}D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}=o_{P}(n^{-1/2})$, for $v=1,\ldots,V$. One could also carry out a pooled-TMLE-update step by minimizing the cross-validated empirical risk $\epsilon\rightarrow 1/V\sum_{v}P_{n,v}^{1}L(P_{n,v}^{0}(\epsilon))$ using the same least favorable path $\{P_{n,v}^{0}(\epsilon):\epsilon\}$. In both cases, we end up with $P_{n,v}^{\star}$, $v=1,\ldots,V$, that solves $1/V\sum_{v=1}^{V}P_{n,v}^{1}D_{{\cal M}_{n,w,v},P_{n,v}^{\star}}=o_{P}(n^{-1/2})$. We note that if ${\cal M}_{w,n,v}$ is a parametric model, and a sample size of $n/V$ is large enough to train the parameters of this working model, then one could estimate $P_{0}$ with an MLE $P_{n,v}^{\star}=\arg\min_{P\in{\cal M}_{w,n,v}}P_{n,v}^{1}L(P)$, thus not using an initial estimator. Using a least favorable path with many more extra parameters in the pooled-TMLE allows one to use a simple initial estimator (less data-adaptive), the pooled-TMLE starts resembling a parametric MLE over the cross-validated empirical risk.
@@
$\psi_{n}^{\star}\equiv\frac{1}{V}\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n,v}}(P_{n,v}^{\star}).$
@@
$\psi_{n}^{\star}-\frac{1}{V}\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n,v}}(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{\star},P_{0})+o_{P}(n^{-1/2}).$
We assume the analogue of $\Psi_{{\cal M}_{w}}(P_{0})-\Psi(P_{0})=o_{P}(n^{-1/2})$ as presented in in Section 6, which is now given by
@@
$\psi_{n}^{\star}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{\star},P_{0})+o_{P}(n^{-1/2}).$
The analogue rate of convergence condition is given by $1/V\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{\star},P_{0})=o_{P}(n^{-1/2})$. If conditional on the training sample $P_{n,v}$, $D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}$ falls in a $P_{0}$-Donsker class (note that only the targeting step depends on $P_{n,v}^{1}$, making this a very weak condition), then it follows
$\psi_{n}^{\star}-\Psi(P_{0})=O_{P}(n^{-1/2}).$
Suppose $\parallel D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}-D_{{\cal M}_{w,n,v},P_{n,v,0}^{\star}}\parallel_{P_{0}}=o_{P}(1)$, where $P_{n,v,0}^{\star}$ is the TMLE-update of $P_{n,v}^{0}$ under $P_{0}$ (i.e., maximize $P_{0}L(P_{n,v,\epsilon}^{0})$ instead of $P_{n,v}^{1}L(P_{n,v,\epsilon}^{0})$ for a universal least favorable path through $P_{n,v}^{0}$). Then, we obtain
$\psi_{n}^{\star}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v,0}^{\star}}+o_{P}(n^{-1/2}).$
For each $v$, we have that $(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v,0}^{\star}}$ is a sum of mean zero independent random variables so that this term will be asymptotically normal if the variance converges to a fixed $\sigma^{2}_{0}$. Let’s assume that $\parallel D_{{\cal M}_{w,n,v},P_{n,v,0}^{\star}}-D_{{\cal M}_{w,n,v},P_{0}}\parallel_{P_{0}}=o_{P}(1)$ so that we obtain
$\psi_{n}^{\star}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{0}}+o_{P}(n^{-1/2}).$
@@
$\sigma_{w,n}^{-1}(\psi_{n}^{\star}-\psi_{0})=\sigma_{w,n}^{-1}\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{0}}/\sigma_{w,n,v}+o_{P}(n^{-1/2}).$
@@
Assume $1/V\sum_{v}P_{n,v}^{1}D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}=o_{P}(n^{-1/2})$. We have
$\psi_{n}^{\star}-\frac{1}{V}\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n,v}}(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{\star},P_{0})+o_{P}(n^{-1/2}).$
@@
$\psi_{n}^{\star}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{\star},P_{0})+o_{P}(n^{-1/2}).$
Rate of convergence condition: Assume $1/V\sum_{v=1}^{V}R_{{\cal M}_{w,n,v}}(P_{n,v}^{\star},P_{0})=o_{P}(n^{-1/2})$.
Weak Donsker class condition: Assume, conditional on the training sample $P_{n,v}$, $D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}$ falls in a $P_{0}$-Donsker class (note that only the targeting step depends on $P_{n,v}^{1}$, making this a very weak condition). Then,
$\psi_{n}^{\star}-\Psi(P_{0})=O_{P}(n^{-1/2}).$
Consistency of TMLE $P_{n,v}^{\star}$ to $P_{0}$: Assume that $\parallel D_{{\cal M}_{w,n,v},P_{n,v,0}^{\star}}-D_{{\cal M}_{w,n,v},P_{0}}\parallel_{P_{0}}=o_{P}(1)$. Let $\sigma_{w,n,v}^{2}=P_{0}D_{{\cal M}_{w,n,v},P_{0}}^{2}$ and $\sigma_{w,n}=1/V\sum_{v=1}^{V}\sigma_{w,n,v}$. Then,
$\sigma_{w,n}^{-1}(\psi_{n}^{\star}-\psi_{0})=\sigma_{w,n}^{-1}\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n,v},P_{0}}/\sigma_{w,n,v}+o_{P}(n^{-1/2}).$
@@
Meta-HAL-MLE to generate working model for $\tau_{0}$: To obtain extra signal in the data for the second regression targeting $\tau_{0}$ we might apply cross-fitting again, analogue to above for the parametric regression working model. Let $P_{n,v}$ and $P_{n,v}^{1}$ be the training and validation sample for the $v$-th sample split, respectively, and $\theta_{n,v}=\hat{\Theta}(P_{n,v})$ be the resulting training sample fit of $\theta_{0}$, $v=1,\ldots,V$. We can define $\beta_{n}^{C}=\arg\min_{\beta,\parallel\beta\parallel\leq C}1/V\sum_{v=1}^{V}P_{n,v}^{1}L(\theta_{n,v}+\sum_{j}\beta_{j}S\phi_{j})$ as the minimizer of the cross-fitted empirical risk. We can select $C$ with cross-validation resulting in a $\beta_{n}^{cf}$ which then generates the working model $\{\sum_{j,\beta_{n}^{cf}(j)\not=0}S\beta_{j}\phi_{j}:\beta\}$ for $\tau_{0}$. One can think of $\theta_{n}+\sum_{j}\beta_{n}^{cf}(j)S\phi_{j}$ as a meta-HAL-MLE, where cross-validation selects the best $\beta$-specific estimator $\hat{\theta}(P_{n})+\sum_{j}S\beta_{j}\phi_{j}$, under an $L_{1}$-norm constraint on $\beta$. Carrying out this meta-HAL-MLE (using cross-validation to select $C$) implies the semiparametric working model ${\cal Q}_{w,sp,n}=\{\theta+\sum_{j,\beta_{n}^{cf}(j)\not=0}\beta_{j}S\phi_{j}:\beta,\theta\}$.
@@
Imagine that one has access to a previous related study that collected the same covariates and outcome with possibly a different treatment. One could use this previous study to learn a working model for the outcome regression $\bar{Q}_{0}$, possibly apply some undersmoothing to make it not too adaptive towards the true regression function in that previous study. One wants to make sure that that study has a sample size larger or equal than the one in the current study so that the resulting working model is flexible enough for the sample size of the current study. One could now use this ${\cal M}_{w}$ as working model in the A-TMLE. One can now either use a TMLE or CV-TMLE of the target parameter $\Psi_{{\cal M}_{w}}(P_{0})$. Below we present the theorem for the resulting adaptive CV-TMLE. The advantage of learning the working model on an external data set is that it allows us to establish asymptotic normality without requiring that $D_{{\cal M}_{w},P_{0}}$ converges to a fixed $D_{{\cal M}_{0},P_{0}}$. One still needs that ${\cal M}_{w}$ approximates $P_{0}$ in the sense that $\Psi_{{\cal M}_{w}}(P_{0})-\Psi(P_{0})=o_{P}(n^{-1/2})$, but as we argued in [adml] that condition can be achieved without relying on ${\cal M}_{w}$ to converge to a fixed oracle model ${\cal M}_{0}$ (in essence only relying on ${\cal M}_{w}$ to approximately capture $P_{0}$).